OpenStudy (anonymous):

can anyone help me with this question Use the fact that f(-2)=0 to factor f(x)=x^3-3x^2-16x-12

OpenStudy (caozeyuan):

well, if f(-2)=0, x+2 has to be a factor right?

OpenStudy (anonymous):

oh yeah i never realized that thanks for pointing that out for me

OpenStudy (caozeyuan):

the other factor has one degree lower than the given cuz x+2 is first order

OpenStudy (caozeyuan):

which means the factor is (x+2)(ax^2+bx+c) and a is not 0

OpenStudy (caozeyuan):

now set two equal to each other, expand LHS and compare coefficients to determine a,b,c if the secondary polynomial has real roots, rewrite the polynomial as (x+x1)(x+x2) where x1and x2 are the roots

OpenStudy (anonymous):

im lost now

OpenStudy (caozeyuan):

Ok lets work this out, you undertand that x+2 is one of the factors right

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

and the other part has the form ax^2+bx+c,right?

OpenStudy (anonymous):

where did you get the ax and the bx and the c

OpenStudy (caozeyuan):

Do you mean how did I come up with this form?

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

well, 5^3 can be factored into 5^2 and 5, right

OpenStudy (caozeyuan):

same thing here, we pull a x out of x^3,the left part has to be x^2

OpenStudy (anonymous):

ok

OpenStudy (caozeyuan):

so we have the answer already, it is (x+2)(ax^2+bx+c)

OpenStudy (caozeyuan):

therefore it has to equal to x^3-3x^2-16x-12

OpenStudy (anonymous):

so whats the answer

OpenStudy (caozeyuan):

we are getting there, still a long way to go actually

OpenStudy (caozeyuan):

so we have (x+2)(ax^2+bx+c)=x^3-3x^2-16x-12

OpenStudy (caozeyuan):

therefore ax^3+(2a+b)x^2+(2b+c)x+2c=RHS

OpenStudy (caozeyuan):

now we have a=1,2c=-12,2a+b=-3, so you can solve for a b and c

OpenStudy (anonymous):

ok

OpenStudy (caozeyuan):

so tell me what are a b and c

OpenStudy (anonymous):

b=-5

OpenStudy (caozeyuan):

ok, good

OpenStudy (caozeyuan):

now set ax^2+bx+c=0, is there any real root?

OpenStudy (anonymous):

no

OpenStudy (caozeyuan):

well, yes actually b^2-4ac text remember?

OpenStudy (anonymous):

yeah sorry

OpenStudy (caozeyuan):

so (-5)^2-4*-6=49, which is neat, because it is 7^2

OpenStudy (caozeyuan):

now can you solve the equation?

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

so what is the solution?

OpenStudy (anonymous):

isn't 49

OpenStudy (caozeyuan):

I am the quadratic equation what are the roots, there are two, if you haven't realized already

OpenStudy (anonymous):

well thanks for the help i got to go now